OpenStudy (anonymous):

how do you factor 25m^3

jimthompson5910 (jim_thompson5910):

25 = 5*5 and m^3 = m*m*m So 25m^3 = 5*5*m*m*m

OpenStudy (anonymous):

i need it in simplest form

jimthompson5910 (jim_thompson5910):

That's as simple as it gets since every factor is a prime factor

jimthompson5910 (jim_thompson5910):

oh I guess you can write 5*5 as 5^2, but that's not much of a difference